Radio stations in the U.S. have call letters that begin with K or W. Some have a total of 3 letters; others have a total of 4 letters.
(a) How many sets of 3-letter call letters are possible if:

(i) the last two letters can be repeated?




(ii) the last two letters cannot be repeated?




(b) How many sets of 4-letter call letters are possible if:
(i) the last three letters can be repeated?




(ii) the last three letters cannot be repeated?

Answer by vheroli(126)

a)
i. 2*26*26=1352
ii. 2*25*24=1200
b)
i. 2*26*26*26=35152
ii. 2*25*24*23=27600